The interior angles of a polygon equal 180°. How many sides does this polygon have?
Found 2 solutions by stanbon, jeseca_1964:
Answer by stanbon(75887)

The interior angles of a polygon equal 180°. How many sides does this polygon have?
-----------
Every regular polygon with n sides has interior angles
that add up to (n-2)90 degrees
--------------------------------------
Equation:
(n-2)90 = 180
n-2 = 2
n = 4 sides
=======================
Cheers,
Stan H.

Answer by jeseca_1964(5)

180(n-2) = 180
180(n-2)/180 = 180/180
n-2 = 1
n = 1 + 2
n = 3